a river 4 miles wide is following at rate of 2 miles/hr . The minimum time taken by a boat to cross the river with a speed v = 4 miles / hour (in still water ) is approximately  
(a) 1 hr 9 min
(b) 2 hr 7 min
(C) 1 hr 12 min
(d) 2 hr 25 min
 

Grade:9

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The net speed of the boat would be given by:
v_{b}=\sqrt{4^{2}+2^{2}}=\sqrt{16+4}=\sqrt{20}=2\sqrt{5} m/s
The time it would take would be equal to
t=\frac{4}{4} s=1 hr
Regards
Sumit